# StableSketcher: Enhancing Diffusion Model for Pixel-based Sketch Generation via Visual Question Answering Feedback

### v1

@@ 3.2.2 Negative Sketch Data Construction
To build a high-quality negative sample set for contrastive learning, we constructed 1,693 sketch–caption–QA triplets. All images were generated using Stable Diffusion v2.1, and captions were produced via GPT-4o and subsequently refined through human post-editing. QA pairs were generated using the TIFA framework [10], which combines LLaMA 2 [30] for question generation and UnifiedQA [12] for answer validation. Negative samples were generated through a systematic procedure focused on producing images that diverge from the target sketch style. The prompts were explicitly designed to generate characteristics inconsistent with the black-on-white sketch aesthetic, including detailed line drawings, colored elements, textured or non-white backgrounds, and heavy shading. Following generation, all candidate images underwent human filtering to ensure inclusion criteria and semantic fidelity.
@@ 3.3 Sketch-Caption Pair
To construct high-quality sketch–caption pairs, we initially experimented with several state-of-the-art image captioning models. Although models like BLIP-2 Flan T5-xl [18] and mPLUG [17] showed strong performance on general datasets, they often produced generic, incorrect, or overly simplistic captions for sketches. For example, captions such as “a black drawing of a cat” or nonsensical outputs highlighted their inability to capture fine-grained sketch details.
Prompt engineering strategies such as “include three characteristics” and “focus on the eyes” were explored, but they failed to meaningfully improve caption quality. This limitation stems from both the models’ smaller size and their training on object-centric datasets like COCO Captions, which emphasize inter-object relationships over instance-specific descriptions.
To overcome this limitation, larger vision-language models were adopted, namely mPLUG-Owl3 [36] and GPT-4o [11]. These models demonstrated significantly improved ability to generate rich, instance-specific, and semantically faithful captions. For instance, GPT-4o could describe “a playful drawing of a fish with stripes, a small round eye, and a triangular tail fin,” better aligning with the level of abstraction and detail present in our sketches.
All model-generated captions were then refined through a rigorous human correction process to ensure accuracy, especially in counting and spatial relationships [10].
@@ 4 StableSketcher
4.1 Preliminaries
Diffusion Models
@@ 4.1 Preliminaries
Latent Diffusion Models (LDMs)
Stable Diffusion [23] adapts this framework into a latent space for efficiency. Instead of operating directly in pixel space, an image $x_{0}$ is encoded into a latent representation $z$ via a variational autoencoder (VAE), and the diffusion process is carried out in this lower-dimensional space. The training objective then becomes:
@@
Denoising Diffusion Policy Optimization (DDPO)
While diffusion models generate visually realistic images, they may not align closely with input conditions (e.g., textual prompts) or task-specific objectives. Denoising Diffusion Policy Optimization (DDPO)  [1] addresses this limitation by framing the denoising process as a Markov decision process (MDP) [29], where each denoising step is treated as an action. At timestep $t$, the model predicts a denoised sample $x_{t-1}$ and receives a reward $r(x_{t-1},x_{0},y)$ measuring alignment with the conditioning input $y$. The optimization objective is to maximize the expected cumulative reward:
@@
Building upon these foundations, we propose StableSketcher, a training framework that adapts Stable Diffusion for sketch generation while incorporating DDPO with VQA-based feedback and evaluation. As illustrated in Figure 4, StableSketcher accepts as inputs a textual prompt $\mathbf{T}$ and a human-drawn sketch $\mathbf{I}$ from the SketchDUO dataset. First, the VAE is fine-tuned on sketch images to improve reconstruction fidelity and stability. Next, the Stable Diffusion UNet is trained with the adapted VAE to synthesize human-drawn–style sketches conditioned on $\mathbf{T}$. Finally, Stable Diffusion is fine-tuned via DDPO using our VQA-based reward, which parses elements from $\mathbf{T}$ and evaluates them individually to strengthen prompt fidelity across samples. The training procedure comprises two main stages executed sequentially: (i) Stable Diffusion training and (ii) VQA-driven DDPO feedback, detailed in the following subsections.
@@ 4.2 Training Stable Diffusion
VAE Fine-tuning for Sketch Reconstruction
@@
UNet Fine-tuning for Text-Aligned Sketch Generation
We perform UNet fine-tuning on Stable Diffusion using sketch–caption pairs from the SketchDUO dataset to adapt the model for generating human-drawn style sketch images. As illustrated in Figure 4, the frozen VAE is replaced with our enhanced VAE to better capture sketch-specific representations. Text prompts are incorporated into the UNet through a cross-attention mechanism, enabling the model to effectively align the denoising process with the given prompt. Furthermore, the denoising diffusion probabilistic models (DDPM) [8] scheduler is employed to ensure a stable and consistent diffusion process during training. We follow the original noise prediction objective of Stable Diffusion [23] for UNet fine-tuning.
@@ 4.3 VQA-Guided Fine-tuning with DDPO
Design of VQA-Based Reward Function
@@
VQA-Based Reward Function with SketchDUO QA Triplets
@@
For each image, there are $N+M$ QA pairs, consisting of $N$ instance-related questions and $M$ sketch-related questions. The weighting ratio for each component is controlled by $\alpha$, where $0\leq\mathcal{R}_{\text{VQA}}\leq 1$. We set $\alpha=0.5$ in our experiments:
@@
VQA Model Fine-tuning for Accurate Reward Signals
The accuracy of the VQA model directly impacts the reliability of the reward signal and thus the quality of policy updates. High rewards are assigned when generated sketches align with prompt elements, while mismatches yield lower rewards, guiding the policy toward faithful sketch generation. To improve sketch understanding, we fine-tune the VQA model on the SketchDUO QA set, using 80% of the data for training and 20% for evaluation. A comparison of the baseline mPLUG-large and our fine-tuned model is presented in Table 3.
@@ 5.1 Implementation Details
In this study, a dataset consisting of 3K sketch image-caption pairs, was divided into a training set of 1.8K and a test set of 1.2K, with data utilized differentially across learning stages. In the VAE fine-tuning process, 1.8K sketch images were used for training, while the Stable Diffusion UNet learning employed 1.8K sketch image-caption pairs. The DDPO algorithm-based reinforcement learning was conducted using 1.8K sketch images and approximately 29K corresponding QA pairs. In the process of fine-tuning the mPLUG-large VQA model, additional negative samples were integrated. These comprised about 1.7K sketch images in undesired styles, their corresponding captions, and approximately 17K question-and-answer sets.
Baseline
@@
This paper employs Stable Diffusion v1.5 [23] as the baseline model due to its balanced performance in both image quality and text-image alignment compared to other versions.
Evaluation metrics
@@
To evaluate the quality of the generated images, we adopt five metrics that encompass both image quality and text-image alignment. For image quality assessment, we employ Fréchet Inception Distance (FID) [7] and LPIPS [38]. For text-image alignment, we leverage CLIPScore [6], BERTScore [40], and TIFAScore [10].
@@
Table 4: Quantitative evaluation of generated images using FID, CLIPScore, BertScore, and TIFAScore metrics for different configurations of Stable Diffusion models. The “+ Fine-tuning” rows indicate that fine-tuning was applied to the corresponding base model, while “+ VAE fine-tuning” rows represent the additional application of VAE fine-tuning on top of the fine-tuned model.
Method FID $\downarrow$ CLIPScore $\uparrow$ BertScore $\uparrow$ TIFAScore $\uparrow$
Stable Diffusion v1.5 207.59 $\pm$ 22.3 34.00 $\pm$ 2.6 0.89 $\pm$ 0.03 0.59 $\pm$ 0.15
   + UNet fine-tuning 161.94 $\pm$ 20.3 36.05 $\pm$ 2.6 0.89 $\pm$ 0.03 0.68 $\pm$ 0.13
    + VAE fine-tuning 143.68 $\pm$ 16.6 35.48 $\pm$ 2.5 0.88 $\pm$ 0.03 0.68 $\pm$ 0.12
Stable Diffusion v2.1 230.78 $\pm$ 22.7 31.13 $\pm$ 3.4 0.88 $\pm$ 0.03 0.53 $\pm$ 0.15
   + UNet fine-tuning 144.46 $\pm$ 25.7 34.79 $\pm$ 2.7 0.88 $\pm$ 0.03 0.67 $\pm$ 0.13
    + VAE fine-tuning 172.35 $\pm$ 14.5 34.11 $\pm$ 2.8 0.88 $\pm$ 0.03 0.65 $\pm$ 0.13
Preliminary experiments
Initially, we evaluated mPLUG-large on the SketchDUO QA test sets, achieving an accuracy of $61.3\%$, as shown in Table 3. Given the insufficient accuracy of this model, we proceeded to fine-tune mPLUG-large on SketchDUO. The fine-tuned model reached an accuracy of $89.3\%$, significantly improving its performance.
@@ 5.2 Quantitative results
Based on Table 4, Stable Diffusion v1.5 demonstrated superior baseline performance compared to v2.1. While both models showed improvements with UNet fine-tuning, v1.5 achieved greater enhancements in reducing FID and increasing TIFAScore. With additional VAE fine-tuning, v1.5 recorded the lowest FID of $143.68$ and the highest TIFAScore of $0.68$, delivering the best overall results. In contrast, VAE fine-tuning had minimal impact on v2.1’s performance. Therefore, Stable Diffusion v1.5 with UNet and VAE fine-tuning, which offers the best performance in text-image alignment and image quality, was selected for use with the DDPO algorithm.
@@
Figure 7 illustrates the training process of the DDPO algorithm with the proposed reward function. For the text prompts $T_{1}$ and $T_{2}$, the generated images progressively aligned better with the prompts as training progressed. The reward for $T_{1}$ increased from $0.77$ to $0.85$, while the reward for $T_{2}$ improved from $0.59$ to $0.82$. For $T_{1}$, the reward initially increased rapidly and then stabilized, with the generated images progressively reflecting the finer details of the prompt. In contrast, $T_{2}$ showed a steady improvement in the reward function throughout training, and the corresponding generated images consistently aligned more closely with the prompt. Meanwhile, for prompt $T_{3}$, the reward remained nearly unchanged throughout the training process, indicating limitations in achieving full prompt fidelity in the generated images. The right graph of Figure 7 visualizes the reward progression over the training steps, demonstrating that the proposed reward function effectively enhances text-image alignment and stabilizes the learning process. The DDPO algorithm consistently generates images with higher prompt fidelity as training progresses, validating the effectiveness of the proposed reward function.
5.3 Qualitative results.
@@
Figure 8 compares image quality across Stable Diffusion variants and our framework, StableSketcher. Stable Diffusion v1.5 and v2.1 show characteristic failures—v1.5 often over-details or drifts from the text, while v2.1 is more abstract and inconsistent. Fine-tuning the UNet improves simplicity and prompt alignment; tuning both UNet and VAE further boosts fidelity but remains unstable for the “white background” and fully accurate instance generation. In contrast, the proposed StableSketcher applies DDPO, an RL-based policy optimization algorithm, to overcome the limitations of baseline models and achieve the best results. The images generated by StableSketcher resemble human-drawn sketches and faithfully reflect the detailed conditions of the text prompts. Additionally, the results from StableSketcher were the most similar to the ground truth images, demonstrating the effectiveness of the proposed framework. Comprehensive qualitative results for all classes, together with the corresponding text prompts, are provided in the Appendix.

### v2

@@ 3.2.2 Negative Sketch Data Construction
To build a high-quality negative sample set for contrastive learning, we constructed 1,693 sketch–caption–QA triplets. All images were generated using Stable Diffusion v2.1, and captions were produced with GPT-4o using a separate prompt tailored to off-style sketches. Unlike the positive template, this prompt explicitly allowed descriptions of non-white backgrounds, color, excessive detail, and shading, which are central attributes of the negative subset. The generated captions were then manually revised to preserve semantic correctness while retaining the undesired stylistic traits needed for contrastive supervision. The exact prompt is provided in Table 2.
QA pairs were generated using the TIFA framework [16], which combines LLaMA 2 [45] for question generation and UnifiedQA [19] for answer validation. Negative samples were generated through a systematic procedure focused on producing images that diverge from the target sketch style. These controlled stylistic deviations serve as contrastive supervision, enabling the model to better distinguish desired sketch attributes from undesired ones. The prompts were explicitly designed to generate characteristics inconsistent with the black-on-white sketch aesthetic, including detailed line drawings, colored elements, textured or non-white backgrounds, and heavy shading. Following generation, all candidate images underwent human filtering to ensure inclusion criteria and semantic fidelity.
@@ 3.3 Sketch-Caption Pair
To construct high-quality sketch–caption pairs, we initially evaluated several image captioning models, including BLIP-2 Flan T5-xl [26], mPLUG [25], mPLUG-Owl3 [53], and GPT-4o [18]. Smaller captioning models often produced generic, incorrect, or overly simplified descriptions for sketches, especially for instance-specific attributes such as counts, local shapes, and background conditions. We therefore adopted GPT-4o as the primary captioning model and used template-constrained prompts tailored to SketchDUO. This caption construction process followed a human-in-the-loop pipeline, where GPT-4o first generated candidate captions and human annotators then verified and corrected them before QA generation.
As summarized in Table 2, the positive prompt constrained the description to a black-on-white single-object sketch format, whereas the negative prompt explicitly allowed off-style properties such as non-white backgrounds, color, and heavy shading. This prompt design was intended to reflect the distinct roles of the two subsets in the contrastive dataset construction pipeline.
All GPT-4o captions were then manually revised before QA generation. Human post-editing focused on correcting object identity, counts, local attributes, and background descriptions while preserving the intended subset-specific style. For positive sketches, annotators normalized the captions to remain faithful to the black-line, white-background sketch definition. For negative sketches, annotators preserved semantically correct object descriptions while explicitly retaining undesired stylistic traits such as color, excessive detail, non-white background, and shading.
@@ 4 StableSketcher
4.1 Backgrounds
4.1.1 Diffusion Models
@@
4.1.2 Latent Diffusion Models (LDMs)
Stable Diffusion [38] adapts this framework into a latent space for efficiency. Instead of operating directly in pixel space, an image $x_{0}$ is encoded into a latent representation $z$ via a VAE, and the diffusion process is carried out in this lower-dimensional space. The training objective then becomes:
@@ 4.1.2 Latent Diffusion Models (LDMs)
4.1.3 Denoising Diffusion Policy Optimization
While diffusion models generate visually realistic images, they may not align closely with input conditions (e.g., textual prompts) or task-specific objectives. DDPO  [4] addresses this limitation by framing the denoising process as a Markov decision process (MDP) [44], where each denoising step is treated as an action. At timestep $t$, the model predicts a denoised sample $x_{t-1}$ and receives a reward $r(x_{t-1},x_{0},y)$ measuring alignment with the conditioning input $y$. The optimization objective is to maximize the expected cumulative reward:
@@ 4.1.3 Denoising Diffusion Policy Optimization
Building upon these foundations, we propose StableSketcher, a training framework that adapts Stable Diffusion for sketch generation and incorporates DDPO with VQA-based feedback. As illustrated in Figure 5, StableSketcher takes as input a textual prompt $\mathbf{T}$ and a human-drawn sketch $\mathbf{I}$ from the SketchDUO dataset. First, the VAE is fine-tuned on sketch images to enhance reconstruction. Next, the U-Net of Stable Diffusion is trained with the adapted VAE to generate human-drawn style sketches conditioned on $\mathbf{T}$. Finally, Stable Diffusion is fine-tuned with DDPO using our VQA-based reward, which extracts elements from $\mathbf{T}$ and evaluates them individually to improve prompt fidelity. The training procedure can be divided into two main stages: (i) Training Stable Diffusion and (ii) VQA Feedback using DDPO, detailed in the following subsections.
@@ 4.2 Training Stable Diffusion
4.2.1 VAE Fine-tuning for Sketch Reconstruction
@@
4.2.2 U-Net Fine-tuning for Text-Aligned Sketch Generation
We perform U-Net fine-tuning on Stable Diffusion using sketch–caption pairs from the SketchDUO dataset to adapt the model for generating human-drawn style sketch images. As illustrated in Figure 5, the frozen VAE is replaced with our enhanced VAE to better capture sketch-specific representations. Text prompts are incorporated into the U-Net through a cross-attention mechanism, enabling the model to effectively align the denoising process with the given prompt. Furthermore, the denoising diffusion probabilistic models (DDPM) [14] scheduler is employed to ensure a stable and consistent diffusion process during training. We follow the original noise prediction objective of Stable Diffusion [38] for U-Net fine-tuning.
@@ 4.3 VQA-Guided Fine-tuning with DDPO
4.3.1 Design of VQA-Based Reward Function
@@
4.3.2 Implementation VQA-Based Reward Function with SketchDUO QA Triplets
@@
For each image, there are $N+M$ QA pairs, consisting of $N$ instance-related questions and $M$ sketch-style questions. The weighting ratio is controlled by $\alpha$, where $0\leq\mathcal{R}_{\text{VQA}}\leq 1$. We set $\alpha=0.5$ in our experiments:
@@
4.3.3 VQA Model Fine-tuning for Accurate Reward Signals
The accuracy of the VQA model directly impacts the reliability of the reward signal and thus the quality of policy updates. High rewards are assigned when generated sketches align with prompt conditions, whereas mismatches are assigned lower rewards, guiding the policy toward faithful sketch generation. To improve sketch understanding, we fine-tune the VQA model on the SketchDUO QA set, using 80% of the data for training and 20% for evaluation. This fine-tuning uses both positive and negative sketch-QA triplets, allowing the model to learn not only whether a generated image matches the prompt semantics, but also whether it conforms to the intended sketch style. Consequently, the reward function can reflect both desired and undesired stylistic patterns. A comparison of the baseline mPLUG-large model and our fine-tuned model is presented in Table 3.
@@ 5.1.1 Dataset: SketchDUO
In this study, the sketch-image-caption dataset was divided into training and test sets using a 6:4 split, and each subset was used differently across learning stages. For VAE fine-tuning, sketch images from the training split were used for optimization, whereas the Stable Diffusion U-Net was trained on sketch-image-caption pairs from the same split. DDPO-based reinforcement learning was performed using training sketches and their associated QA pairs. In addition, fine-tuning of the mPLUG-large VQA model incorporated negative samples in undesired styles together with their corresponding captions and QA pairs.
5.1.2 Training configuration
Unless otherwise noted, all modules were trained on two NVIDIA RTX A6000 GPUs, and we report the main hyperparameters and approximate wall-clock training times for reproducibility. VAE fine-tuning was conducted for 15 epochs on a single NVIDIA RTX 3060 Ti GPU using AdamW (learning rate $1\times 10^{-5}$, batch size 2, $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay 0.01) with the VGG-based LPIPS and reconstruction losses; this stage took approximately 11 hours. U-Net fine-tuning then used SketchDUO at a resolution of $512\times 512$ with center cropping and random horizontal flipping. We fine-tuned only the U-Net while freezing the VAE and CLIP text encoder, using AdamW (learning rate $1\times 10^{-5}$, batch size 1), gradient accumulation of 1, gradient checkpointing, EMA, and a constant learning-rate scheduler without warmup for 100 epochs; this stage took approximately 30 hours.
DDPO-based fine-tuning started from our VAE-adapted Stable Diffusion checkpoint and optimized LoRA parameters using the TIFAScore-based reward. We trained for 800 epochs with batch size 1, gradient accumulation of 1, learning rate $3\times 10^{-4}$, maximum gradient norm 1, and Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, weight decay $10^{-4}$), with FP16 mixed precision and TF32 enabled. Sampling used 50 denoising steps, guidance scale 5, $\eta=1$, and one sampled batch per epoch; this stage took approximately 16 hours. For VQA fine-tuning, we trained mPLUG-large on the SketchDUO QA set for 6 epochs using both positive and negative sketch–QA triplets, with an 80:20 training/evaluation split; this stage took approximately 36 hours, and the resulting accuracy is reported in Table 3.
5.1.3 Baselines
@@
Based on the results in Figure 8 and Table 4, this paper employs Stable Diffusion v1.5 [38] as the baseline model due to its balanced performance in both image quality and text-image alignment compared to other versions.
5.1.4 Evaluation metrics
@@
To evaluate the quality of the generated images, we adopt five metrics that encompass both image quality and text-image alignment. For image quality assessment, we employ Fréchet Inception Distance (FID) [13] and LPIPS [57]. For text-image alignment, we leverage CLIPScore [12] and BERTScore [58]. While these four metrics are widely established as standard benchmarks for evaluating diffusion-based sketch generation models [48, 15, 46], they are not specifically tailored to sketch abstraction.
We therefore additionally incorporate TIFAScore [16] and complement the quantitative evaluation with a user study. This inclusion allows us to explicitly measure fine-grained prompt fidelity, providing a more precise assessment of semantic alignment than standard metrics alone. Finally, recognizing that these automated metrics may not fully capture the unique characteristics of sketches, we complemented our quantitative evaluation with a user study.
@@
Table 4: Quantitative evaluation of generated images using FID, LPIPS, CLIPScore, BERTScore, and TIFAScore metrics for different configurations of Stable Diffusion models. The “+ Fine-tuning” rows indicate that fine-tuning was applied to the corresponding base model, while “+ VAE fine-tuning” rows represent the additional application of VAE fine-tuning on top of the fine-tuned model.
Method FID $\downarrow$ LPIPS $\downarrow$ CLIPScore $\uparrow$ BERTScore $\uparrow$ TIFAScore $\uparrow$
[…]
5.2 Preliminary experiments
Initially, we evaluated mPLUG-large on the SketchDUO test sets and obtained an accuracy of 61.3%, as reported in Table 3. Since this performance was insufficient for reliable reward modeling, we fine-tuned mPLUG-large on SketchDUO. The fine-tuned model achieved a substantially improved accuracy of 89.3% in the best setting. Furthermore, the model trained on both positive and negative samples consistently outperformed the version trained on positive samples alone, indicating that negative samples provide useful contrastive supervision for distinguishing desired sketch properties from undesired stylistic attributes.
@@ 5.3 Quantitative results
Based on Table 4, Stable Diffusion v1.5 demonstrated superior baseline performance compared to v2.1. While both models showed improvements with U-Net fine-tuning, v1.5 achieved greater enhancements in reducing FID and increasing TIFAScore. With additional VAE fine-tuning, v1.5 recorded the lowest FID of $143.68$ and the highest TIFAScore of $0.68$, delivering the best overall results.
In contrast, VAE fine-tuning had minimal impact on v2.1’s performance. Therefore, Stable Diffusion v1.5 with U-Net and VAE fine-tuning, which offers the best performance in text-image alignment and image quality, was selected for use with the DDPO algorithm.
@@
Figure 7 illustrates the training process of the DDPO algorithm with the proposed reward function. For the text prompts $T_{1}$ and $T_{2}$, the generated images progressively aligned better with the prompts as training progressed. The reward for $T_{1}$ increased from $0.77$ to $0.85$, while the reward for $T_{2}$ improved from $0.59$ to $0.82$. For $T_{1}$, the reward initially increased rapidly and then stabilized, with the generated images progressively reflecting the finer details of the prompt. In contrast, $T_{2}$ showed a steady improvement in the reward function throughout training, and the corresponding generated images consistently aligned more closely with the prompt.
Meanwhile, for prompt $T_{3}$, the reward remained nearly unchanged throughout the training process, indicating limitations in achieving full prompt fidelity in the generated images. The right graph of Figure 7 visualizes the reward progression over the training steps, demonstrating that the proposed reward function effectively enhances text-image alignment and stabilizes the learning process. The DDPO algorithm consistently generates images with higher prompt fidelity as training progresses, validating the effectiveness of the proposed reward function.
5.4 Qualitative results
@@
Figure 8 compares image quality across Stable Diffusion variants and our framework, StableSketcher. Stable Diffusion v1.5 and v2.1 show characteristic failures, as v1.5 often produces overly detailed outputs or drifts from the text prompt, whereas v2.1 tends to generate more abstract yet inconsistent results. Fine-tuning the U-Net improves simplicity and prompt alignment; tuning both U-Net and VAE further boosts fidelity but remains unstable for the “white background” and fully accurate instance generation.
Figure 9: User study interface used for subjective evaluation. The user study was conducted using a Google Forms interface and consisted of three evaluation criteria: (a) sketch characteristics, (b) prompt fidelity, and (c) human-likeness. For each criterion, participants were presented with five generated images and asked to rank them from 1st to 5th according to the corresponding evaluation guideline.
In contrast, the proposed StableSketcher applies DDPO, an RL-based policy optimization algorithm, to overcome the limitations of baseline models and achieve the best results. The images generated by StableSketcher resemble human-drawn sketches and faithfully reflect the detailed conditions of the text prompts. Additionally, the results from StableSketcher were the most similar to the ground truth images, demonstrating the effectiveness of the proposed framework.